OpenStudy (anonymous):

What is the x-intercept and y-intercept of the graph of 3x - 4y = 12? x-intercept = (-4, 0); y-intercept = (0, 3) x-intercept = (4, 0); y-intercept = (0, -3) x-intercept = (0, -3); y-intercept = (4, 0) x-intercept = (0, 3); y-intercept = (-4, 0)

OpenStudy (callisto):

3x - 4y = 12 3x - 4y -12 =0 x-int = -c/a = -[(-12)/(3)] = 4 y-int = -c/b = -[(-12)/(-4)] = -3 therefore x-int = (4,0), y-int = (0, -3)

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome :)